Apply the special theory of relativity to prove that the time interval ( t ) in the


moving reference frame ( S' ) appears to be longer by a factor "\frac{1}{^{\sqrt{1-\frac{v^{2}}{c^{2}}}}}" observer in the stationary frame ( S ) .


1
Expert's answer
2020-09-24T11:08:04-0400

t="\dfrac{t^`}{1-\dfrac{v}{c^`\scriptscriptstyle k^`\Rightarrow k}}" = "\dfrac{t^`}{1-\dfrac{v}{c}}"

whereas, the same perceived time interval in k from the perspective of k.is given by,

"t=t^`(\dfrac{t^`}{1-\dfrac{v}{c^`\scriptscriptstyle k^`\Rightarrow k}})=t^`(1+{V \over C})"

"t={t^` \over 1-{V \over C}}=t^`(1+{V \over C})"

leading to v=0 unless t is transformed by a certain factor say "\gamma" with respect to k, and by other factor say "\beta" with respect to k'. This leads to;

"t={\beta t^` \over 1-{V \over C}}=\gamma t^`(1+{V \over C})"

which can be satisfied only if "\beta =1 / \gamma"

which results to;

"\gamma =" "\cfrac{1}{\sqrt{1 - \cfrac{V^2}{C^2}}}"


"\therefore" "t=\dfrac{t^`}{\gamma(1-\dfrac{v}{c})}" ="\gamma t^`(1+{V \over C})"

hence the prove that;

"\gamma =" "\cfrac{1}{\sqrt{1 - \cfrac{V^2}{C^2}}}"